OpenStudy (anonymous):

2sin^2 theta + sin theta =1

OpenStudy (anonymous):

\[2\sin ^{2} \theta + \sin \theta = 1\] 2sin^2 theta + sin theta -1 =0 or, 2 sin^2 theta + 2 sin theta - sin theta -1 = 0 or, 2sin theta (sin theta +1) - 1 (sin theta +1) = 0 or, (sin theta +1) (2sin theta -1) =0 either, sin theta +1=0 [solve it urself) or, 2sin theta -1=0

OpenStudy (anonymous):

thank you so much!!!!!

OpenStudy (anonymous):

happy to help :)

OpenStudy (anonymous):

:D